# MAVIL: Design of a Multidimensional Assessment of Visual Data Literacy and its Application in a Representative Survey

### v2

@@ 6.2 Limitations
As with any study, certain limitations should be acknowledged. These include aspects related to the surveyed population, such as representation boundaries. Respondents below the age of $18$ and above the age of $74$ were not considered in the survey because of legal age as a consent criterion and due to availability reasons (see Section 4.3.3). The respondents share the commonality of being incentivized by a market research organization, which may introduce a bias toward individuals with frequent online survey participation and, therefore, higher digital skills. However, the approach we have chosen to select respondents is also standard in other representative surveys (e.g., [15, 39, 16]).
We recognize that assessment results are likely to vary based on respondents’ socio-cultural and geographical backgrounds, which limits the scalability and comparability of MAVIL. While it allows for inputs from different perspectives, including optional open-ended responses, MAVIL has not been tested for transferability across other cultural or contextual settings. Instead, we used MAVIL to support context-specific interpretation and demonstrated in our representative survey (see Section 4.3) that age and self-perceived numeracy significantly influenced data reading performance (see Section 5.3.2).
@@
MAVIL’s integration of self-perceived ability ratings offers valuable insights into how people engage with visualizations beyond task-based performance, but it may potentially introduce sources of noise. Self-assessments can be influenced by factors like over- or underestimation, social desirability, or ambiguity in question interpretation [36, 57]. While we have piloted MAVIL in several qualitative rounds to ensure clarity and ease of interpretation (see Section 4.2), biases inherent to self-assessment may remain. Nevertheless, as discussed in the Background (see Section 2.3), self-perceived ability ratings are a recognized approach in the study of abilities and continue to provide valuable insights [77, 36, 50].
The visualizations used in the survey, a line chart and a bar chart on climate data, were chosen for their visual simplicity and assumed broad familiarity (see Section 4.3.2). This helped ensure accessibility to general audiences but may limit the scope of our insights. Two visualizations cannot be representative of the variety, complexity levels, and design possibilities of data visualizations. MAVIL should be tested with more diverse visualizations, such as interactive visualizations or intentionally misleading ones, to evaluate how it performs under varying complexity and design conditions.
@@
We deliberately selected climate visualizations for their real-world relevance, aiming to assess how people engage with meaningful, high-stakes information rather than artificial stimuli (see Section 4.1.1). While such topics may evoke emotionally charged or politically influenced reactions, we argue that these responses are a valid part of the interpretation process and reflect how visualizations are encountered in practice. Studying potentially polarizing visuals is therefore, not only justified but essential for understanding real-world information retrieval and addressing misinformation [55, 26]. To account for possible bias, we let respondents assess their topic interest (see MAVIL Q6a in Figure 1). We also examined unsolicited climate-related commentary in our qualitative data and tested whether these participants differed in their data reading performance, finding no significant effect (see Section 5.4.2). That said, we acknowledge that topic-related influences, whether related to climate or other subjects, would benefit from further investigation. This could mean including additional questions on topic interest to provide more nuanced contextualization of responses in future surveys.
@@ 6.3 Future Work
Future implementations of MAVIL should be expanded to include a wider variety of visualization formats. Beyond static charts, interactive visualizations are increasingly common in digital communication, possibly requiring different skills [45] and forms of audience engagement [29]. Testing MAVIL on such formats may warrant an additional dimension to assess interaction-based ability facets. Other common visualization types, including infographics, could be explored to reflect broad visual data communication practices [51, 44, 37, 38].
MAVIL could further benefit from integration with existing tools like Mini-VLAT [67], as a combined assessment would allow for further comparison of self-assessed and performance-based results within the same sample. Initial correlations we observed, such as between self-assessed numeracy and data reading performance, highlight promising grounds for such work. While this study was focused on climate data, MAVIL could be adapted to other domains by rephrasing topic specificity in the Contextualizing dimension, offering a promising direction for future work on assessing its transferability.
Finally, MAVIL could incorporate an educative feedback component. Respondents frequently expressed curiosity about their performance in the data reading tasks during qualitative testing. Providing immediate feedback, correct answers, or even a summary of self-assessed literacy strengths and weaknesses could enhance the assessment’s value, encourage reflection, while serving literacy-building [17], and could function as an incentive for respondents.
@@ 7 Conclusion
With MAVIL we contribute a novel approach to assessing Visual Data Literacy by conceptualizing it as a multidimensional ability composed of distinct facets that can be partially self-assessed. Grounded in a theory-driven approach, it introduces six ability dimensions and combines task-based, self-assessed, and open-ended question formats. This expands existing literacy assessments by capturing not only performance measures but also self-perceived ability ratings. We tested MAVIL in what is, to our knowledge, the first representative survey to assess Visual Data Literacy using a multidimensional approach. Validation was supported through a nationally representative sample covering age groups $18$ to $74$ and a balanced male–female gender split. Our MAVIL survey provided nuanced insights into respondents’ Visual Data Literacy and their understanding of two climate data visualizations. The findings revealed self-assessed deficits, including perceived challenges in interpreting simple measurement units and some unfamiliarity with presumably simple chart types. Overall, MAVIL assesses a combination of different yet central aspects of visualization understanding and aims to encourage more nuanced and critical approaches to evaluating Visual Data Literacy in general audiences.

### v3

@@ 6.2 Limitations
The integration of self-perceived ability ratings in MdamV offers valuable insights into how people engage with visualizations beyond task-based performance, but it may potentially introduce sources of noise. Self-assessments can be influenced by factors like over- or underestimation, social desirability, or ambiguity in question interpretation [38, 63]. While we have conducted pilots in several qualitative rounds to ensure clarity and ease of interpretation (see Sec. 4.2), biases inherent to self-assessment may remain. Nevertheless, as discussed in the Background (see Sec. 2.4), self-perceived ability ratings are a recognized approach in the study of abilities and continue to provide valuable insights [84, 38, 54].
Regarding our survey conduction, there are some representation boundaries. Respondents below the age of $18$ and above the age of $74$ were not considered in the survey because of legal age as a consent criterion and due to availability reasons (see Sec. 4.3.3). The respondents share the commonality of being incentivized by a market research organization or related platforms, which may introduce a bias toward individuals with frequent online survey participation and, therefore, higher digital skills. However, the approach we have chosen to select respondents is also standard in other representative surveys (e.g., [16, 41, 17]). We recognize that respondents’ processes of understanding are likely to vary based on their socio-cultural and geographical backgrounds, limiting the scalability and comparability of our results. As emphasized throughout, MdamV was tested with a specific panel and can only be interpreted in relation to the visualizations used.
Since we investigated situated visualization understanding, our survey data, and therefore our results, are tied to the tested visualizations. We selected a static line chart and bar chart for their assumed broad familiarity (see Sec. 4.1.1) to keep the assessment relevant to general audiences. Still, two chart types cannot represent the variety, complexity levels, and design possibilities of data visualizations. Despite this focus, MdamV enabled us to capture and discuss meaningful facets of visualization understanding that may remain hidden even in assessments with more diverse chart sets, such as the Mini-VLAT [73], which relies on a single item per chart type.
@@
We deliberately selected climate visualizations for their real-world relevance, aiming to assess how people engage with meaningful, high-stakes information rather than artificial stimuli (see Sec. 4.1.1). While such topics may evoke emotionally or politically charged reactions, we argue that these responses are a valid part of the understanding process and reflect how visualizations are encountered in practice. Studying potentially polarizing visuals is, therefore, not only justified but essential for understanding real-world information retrieval and addressing misinformation [61, 27]. To account for possible bias, we let respondents assess their topic interest (see Q6a in Fig. 1), just as we examined unsolicited climate-related commentary in our qualitative data and tested whether these participants differed in their data reading performance, finding no significant effect (see Sec. 5.4.2). That said, we acknowledge that topic-related influences, whether related to climate or other subjects, would benefit from further development. This could mean including additional questions on topic interest to provide more nuanced contextualization of responses in future surveys.
@@ 6.3 Future Work
Future applications of MdamV should be expanded to include a wider variety of visualization formats. For example, beyond static charts, interactive visualizations are increasingly common in online communication. They are particularly relevant for further developing our assessment method, since they may require abilities distinct to digital interaction that shape audience engagement [49, 31]. Other common visualization types, including infographics, could be explored to reflect broad visual data communication practices [57, 47, 39, 40].
MdamV could further benefit from integration with existing approaches like Mini-VLAT [73], as a combined assessment would allow for further comparison of self-assessed and task-based results within the same sample. Initial correlations, such as between self-assessed numeracy and data reading performance (see Sec. 5.3.2), highlighted promising grounds for such work. For the evaluation of our method, we focused on visualized climate data. However, MdamV could be adapted to other domains by rephrasing topic specificity in the Contextualizing dimension, offering a promising direction for future work on assessing its transferability.
Another direction is to make the approach more generalizable by adapting it to study setups that address different phenomena in visualization research. This is particularly relevant where user studies have been critiqued for focusing too narrowly on performance measures [28], lacking qualitative insights [48], and context-related factors [42]. For example, a "MdamT" could be developed to assess the situated understanding of visualization tools by examining how users, e.g., perceive their ability to work with the tool and relate to its features. Similarly, in the context of XAI, "MdamX" could offer an approach to how users, e.g., understand and interpret specific explanations for AI systems. Such adaptations may yield insights that directly inform tool or system design and refinement, while remaining compatible with task accuracy measures. The dimensions we propose in our study provide concrete entry points for such user studies, where evaluation through a multidimensional assessment method for understanding would be both valid and worthwhile, with potential to inform other research contexts related to visualization and data understanding.
Finally, MdamV could incorporate an educative feedback component. Respondents frequently expressed curiosity about their performance in the data reading tasks during qualitative testing. Providing immediate feedback, correct answers and explanations could enhance the assessment’s value, encourage reflection, while serving literacy-building [18], and could function as an incentive for respondents.
@@ 7 Conclusion
With MdamV we contribute a method for assessing situated visualization understanding, allowing a range of abilities related to visual data literacy to be evaluated in direct relation to particular visualizations. By combining task-based measures, self-perceived ability ratings, and open-ended question formats, MdamV extends beyond the traditional focus on performance measures. Grounded in a theory-driven approach, MdamV comprises six distinct dimensions of visualization understanding, in reference to learning sciences frameworks. Validation was supported through a representative survey for Austria’s population (ages 18-74, male/female split). This sample provided nuanced insights into the visualization understanding of a general audience, using two climate data visualizations, and revealed self-reported deficits such as perceived difficulties with simple measurement units and unfamiliarity with common chart types. Overall, the evaluation of MdamV suggests that its dimensions provide a valuable combination for assessing facets of visualization understanding. Importantly, MdamV points toward more nuanced and critical methodological approaches for assessing how general audiences interpret data visualizations. This, in turn, advances our understanding of how visualizations can be meaningfully designed and used in public contexts.